When Kenny's age was twice Bala's age,Mike's age was 30.
When Mike's age was twice Kenny's age,Bala's age was 21.
Given that Mike was the oldest among the three persons,
find Bala's age when Mike is 62 years old.

Found 2 solutions by addingup, greenestamps:
Answer by addingup(3677)

Let Kenny be K, Bala B, and Mike M. And since their ages are unknown, let c be the constant.
.
1)
B = 1c
K = 2c
M = 30
.
Let's find their constant difference.
Difference in age between K and Bala = 1c
Difference in age between Mike and K = 30-2c
.
2)
B = 21
K = 21+1c (from above, we know the constant difference between B & K = 1c)
M = 2(21+1c) = 42 + 2c (Mike is twice of K)
Difference between M & K = 42+2c - (21+1c) = 21+1c
.
30-2c = 21+1c
30-3c = 21
3c = 30 - 21 = 9
1c = 3
.
So what's B's age when M = 62? Well, first we have to find the constant difference between B and M.
M - B = 30 - 3 = 27
So if M = 62, B = 62-27 = 35 This is your answer, when Mike is 62 Bala is 35



Answer by greenestamps(13200)



At one point in time, their ages were: M = 30; B = x; K = 2x. (Kenny was twice Bala's age when Mike was 30).

At another point in time, their ages were: M = 2y; B = 21; K = y. (Mike was twice Kenny's age when Bala was 21).

The number of years between those two times is the same for each of them:
30-2y = x-21 = 2x-y
30-2y = x-21 --> x+2y = 51; x-21 = 2x-y --> y-x = 21
x+2y = 51 and y-x = 21 --> 3y = 72 --> y = 24

So Mike was 48 and Kenny was 24 when Bala was 21.

So the difference between Mike's age and Bala's age is 27 years.

So when Mike is 62, Bala will be 35.